# PoseWatch: A Transformer-based Architecture for Human-centric Video Anomaly Detection Using Spatio-temporal Pose Tokenization

### v1

@@ 5.1 Comparison With Pose-based Approaches
PoseWatch-C 85.10 86.70 85.90
PoseWatch-F 83.19 83.70 83.45
PoseWatch-H 85.75 87.23 86.49
To further illustrate the effectiveness of the proposed model, two examples from SHT [35] and CHAD [13] are shown in Figure 5 and Figure 6. The y-axis represents the output anomaly score of PoseWatch-H for all frames of the testing clip. As depicted in Figure 5, PoseWatch-H accurately detects anomalous behavior, maintaining a steady score on normal frames and showing an increased score on anomalous ones. Figure 6 similarly demonstrates that PoseWatch-H produces higher scores for anomalous frames, enabling anomaly detection. However, more noise is evident in Figure 6, and it is not as accurate as in Figure 5, which is also reflected by the lower AUC-ROC and higher EER observed on the CHAD dataset compared to the SHT dataset showcased in Table 1 and Table 2.
5.2 Comparison With Pixel-based Approaches
This manuscript mainly focuses on pose-based approaches. However, we will further explore an additional comparative analysis between PoseWatch and pixel-based methodologies to have a better understanding of PoseWatch’s capabilities. The datasets common between pose-based and pixel-based approaches are SHT [35] and with one reported instance [63], HR-SHT [40]. Notably, CHAD has not yet been employed in pixel-based studies. Therefore, we further compare PoseWatch with SotA pixel-based algorithms on SHT and HR-SHT.
Table 4: Evaluating AUC-ROC of PoseWatch on SHT [35], HR-SHT [40], and CHAD [13] datasets: A comparative analysis of our design variants with and without incorporating relative motion. The best result of each branch is highlighted in gray.
Relative Movement
SHT
HR-SHT
CHAD
PoseWatch-C 82.97 84.80   57.56
85.10 86.70 66.12
PoseWatch-F 81.80 82.80 58.27
83.19 83.70 66.61
PoseWatch-H 84.20 85.47 57.95
85.75 87.23 67.04
@@ 5.2 Comparison With Pixel-based Approaches
As outlined in Section 1 and Section 2, pixel-based approaches have historically been regarded as more precise than pixel-based approaches. Nonetheless, recent studies indicate a shift in this trend. In Table 3, we present a comparative analysis of PoseWatch against current SotA pixel-based algorithms. PoseWatch-H achieves an average AUC-ROC of $86.49\%$ across the SHT and HR-SHT datasets, underscoring its overall superiority. Notably, PoseWatch-H demonstrates superior performance with AUC-ROC scores of $85.75\%$ and $87.23\%$ on the SHT and HR-SHT datasets, respectively. This superior performance is noteworthy considering that PoseWatch is a pose-based approach. This approach inherently exhibits lower bias and reduced susceptibility to background noise, while simultaneously promoting greater privacy and adhering to ethical standards. Furthermore, despite processing less information than its pixel-based counterparts, PoseWatch-H consistently achieves higher AUC-ROC.
6 Ablation Study
6.1 Impact of Relative Pose
The results detailed in Table 4 empirically validate the effectiveness of using relative pose tokenization, as theoretically outlined in Section 3.1. This empirical evidence complements and reinforces the theory that incorporating global movement benefits anomaly detection. The results indicate a consistent improvement across all three variations of the PoseWatch when the relative pose is integrated. Specifically, regarding AUC-ROC, the PoseWatch-H variant incorporating relative movement demonstrates a notable performance enhancement. On the SHT dataset, for instance, the use of relative movement in PoseWatch-H results in an AUC-ROC increase of $1.55\%$ compared to its counterpart without relative movement. This improvement is even more pronounced in the CHAD, where PoseWatch-H with relative movement achieves a $9.09\%$ higher AUC-ROC than the version without relative pose integration. These findings underscore the significant impact of relative pose tokenization on the model’s results.
Table 5: Evaluating AUC-ROC of PoseWatch on SHT [35], HR-SHT [40], and CHAD [13] datasets: A comparative analysis of our design with different methods of tokenization. The best result of each branch is highlighted in gray.
@@ 6.1 Impact of Relative Pose
PoseWatch-C T-PRP 83.85 85.46 65.59
KS-PRP 83.67 85.15 64.69
FS-PRP 83.45 85.13 62.84
ST-PRP 85.10 86.70 66.12
PoseWatch-F T-PRP 82.08 84.47 66.18
KS-PRP 79.53 80.40 64.95
FS-PRP 82.08 83.21 64.29
ST-PRP 83.19 83.70 66.61
PoseWatch-H T-PRP 84.40 86.06 67.06
KS-PRP 83.63 84.94 65.37
FS-PRP 84.68 86.37 64.28
ST-PRP 85.75 87.23 67.04
Figure 8: Box and Whiskers of various tokenization methods and their effectiveness with regards to AUC-ROC over SHT [35], HR-SHT [40], and CHAD [13] datasets. The plots clearly demonstrate the superiority of the ST-PRP method, which outperforms all other tokenization methodologies across all tested datasets.
6.2 Exploring Diverse Tokenization Strategies
@@
On top of all these tokenization strategies, we also add positional encoding to embed order into input sequences. In Table 5, the performance of various tokenization methods is compared. The ST-PRP tokenization method, detailed in Section 3.1, demonstrates superior performance in most cases compared to other approaches. While the T-PRP tokenization outperforms others in two specific instances, purely spatial tokenizations (KS-PRP and FS-PRP) consistently yield suboptimal results. Figure 8 further supports this observation, clearly demonstrating that across all three benchmark datasets, the PoseWatch variants achieve higher overall AUC-ROC when using ST-PRP tokenization. This suggests that the combination of temporal and spatial attention between tokens uncovers crucial information for analyzing human behavior patterns that neither can detect independently. Consequently, the ST tokenization method, which integrates spatial and temporal information, emerges as the most effective approach.
7 Conclusion
@@
In this paper, we introduced methodologies that pave the way for advanced human-centric Video Anomaly Detection (VAD). The novel proposed Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization method, which serves as a key component for high-level human behavior analysis. Combined with our new Unified Encoder Twin Decoders (UETD) transformer core, the proposed PoseWatch architecture demonstrates superior performance in self-supervised human-centric VAD. Extensive benchmarking against SotA methods confirms PoseWatch’s accuracy and robustness. We hope that our contributions will serve as a foundation for future advancements in the field.
Appendix A Ablation Studies Hyperparameters
@@
This section provides an in-depth exposition of the architectural and training hyperparameters of the ablation studies (Section 6). It is aimed at ensuring reproducibility of the results.
A.1 Relative Pose Ablation Setup and Hyperparameters
@@
To reveal the benefit of incorporating relative movement, we use the best PoseWatch model which includes 12 heads and 4 layers with the feed-forward layer size set to 64. This variant is trained with and without relative movement data using the hyperparameters shown in Table 6. The optimal hyperparameters are chosen using a systematic grid search. In all training instances, we have used the Adam optimizer with a weight decay of $55.0e-$ 5 and trained branches for 30 epochs. For the training, the same strategy is used as in Section 4 in the original manuscript; first, the PoseWatch-C (the unified encoder and CTD decoder) is trained. In the next step, the unified encoder is frozen and the FTD decoder is trained. Since both the SHT and HR-SHT utilize identical videos in their training sets, the hyperparameters for both models remain consistent.
@@
Table 6: The training hyperparameters used for the Relative Pose Ablation Study (Section 6.1) on SHT [35], HR-SHT [40] and CHAD [13] datasets. LR and DR refer to the Learning Rate and Dropout Rate.
@@
LR DR LR DR
PoseWatch-C $32.0e-$ 3 $0.1$ $65.0e-$ 6 $0.1$
$51.0e-$ 5 $0.1$ $32.0e-$ 3 $0.1$
PoseWatch-F $45.0e-$ 4 $0.1$ $33.0e-$ 3 $0.1$
$32.0e-$ 3 $0.1$ $45.0e-$ 4 $0.1$
A.2 Tokeniation Ablation Setup and Hyperparamters
To effectively compare various tokenization setups, we carried out a systematic grid search. This approach was not just to identify the optimal training hyperparameters, but also to determine the best architectural choices. This thorough process ensures that each tokenization setup is evaluated at its highest potential, leading to more accurate and reliable comparison results.
Table 7: The design choices used for the Tokenization Ablation Study (Section 6.2)
@@ A.2 Tokeniation Ablation Setup and Hyperparamters
Table 8: The training hyperparameters used for the Tokenization Ablation Study (Section 6.2) on SHT [35], HR-SHT [40] and CHAD [13] datasets. LR and DR refer to the Learning Rate and Dropout Rate.
@@
CHAD
LR DR LR DR
PoseWatch-C T-PRP $65.0e-$ 6 $0.1$ $33.0e-$ 3 $0.1$
KS-PRP $65.0e-$ 6 $0.1$ $33.0e-$ 3 $0.2$
FS-PRP $51.0e-$ 5 $0.1$ $31.0e-$ 3 $0.1$
ST-PRP $51.0e-$ 5 $0.1$ $32.0e-$ 3 $0.1$
PoseWatch-F T-PRP $33.0e-$ 3 $0.1$ $41.0e-$ 4 $0.1$
KS-PRP $41.0e-$ 4 $0.2$ $41.0e-$ 4 $0.1$
FS-PRP $45.0e-$ 4 $0.1$ $51.0e-$ 5 $0.1$
ST-PRP $32.0e-$ 3 $0.1$ $45.0e-$ 4 $0.1$
Table 7 presents the architectural design parameter choices obtained from the grid search on SHT dataset [35]. These parameters are kept the same for other datasets to ensure a fair comparison. Across all experiments, both the unified encoder and the twin decoders consistently adhere to the parameters detailed in Table 7. This consistency ensures a standardized approach in our experimental setup. On the other hand, Table 8 shows the best learning rate and dropout for training per branch. The number of epochs, optimizer, and weight decay are 30, Adam, and $55.0e-$ 5 respectively.
Regarding the training process, same as other tests, PoseWatch-C is initially trained following the strategy outlined in Section 4.2 in the original manuscript, where both the unified encoder and the CTD decoder are trained together. Subsequently, for PoseWatch-F, the unified encoder is frozen, and then only the FTD decoder undergoes further training. As the training sets for both the SHT and HR-SHT consist of the same videos, the hyperparameters are kept uniform for them.

### v2

@@ V-A Comparison With Pose-based Approaches
SPARTA-C 85.10 62.69 73.90
SPARTA-F 83.19 62.29 72.74
SPARTA-H 85.75 63.48 74.62
To further illustrate the effectiveness of the proposed model, two examples from SHT [29] and CHAD [30] are shown in Figure 4 and Figure 5. The y-axis represents the scaled output anomaly score of SPARTA-H for all frames of the testing clip. As depicted in Figure 4, SPARTA-H accurately detects anomalous behavior, maintaining a steady score on normal frames and showing an increased score on anomalous ones. Figure 5 similarly demonstrates that SPARTA-H produces higher scores for anomalous frames, enabling anomaly detection. However, more noise is evident in Figure 5, and it is not as accurate as in Figure 4, which is also reflected by the lower AUC-ROC and higher EER observed on the CHAD dataset compared to the SHT dataset showcased in Table I and Table II.
V-B Comparison With Pixel-based Approaches
This manuscript mainly focuses on pose-based approaches. However, we will further explore an additional comparative analysis between SPARTA and pixel-based methodologies to have a better understanding of SPARTA’s capabilities. The datasets common between pose-based and pixel-based approaches are SHT [29] and NWPUC [31]. Notably, CHAD [30] has not yet been employed in pixel-based studies. Therefore, we further compare SPARTA with SOTA pixel-based algorithms on SHT and NWPUC.
TABLE IV: Evaluating AUC-ROC on SHT [29], HR-SHT [14], CHAD [30], and NWPUC [31] datasets: A comparative analysis of our design variants with and without incorporating relative motion. The best result of each branch is highlighted in gray.
Relative Movement SHT HR-SHT CHAD NWPUC
SPARTA-C 82.97 84.80 57.56 62.52
85.10 86.70 66.12 62.69
SPARTA-F 81.80 82.80 58.27 62.72
83.19 83.70 66.61 62.29
SPARTA-H 84.20 85.47 57.95 63.41
85.75 87.23 67.04 63.48
@@ V-B Comparison With Pixel-based Approaches
As outlined in Section I and Section II, pixel-based approaches have historically been regarded as more precise than pose-based approaches. Nonetheless, recent studies indicate a shift in this trend. In Table III, we present a comparative analysis of SPARTA against current SOTA pixel-based algorithms. SPARTA-H achieves an average AUC-ROC of $74.62\%$ across the SHT and NWPUC datasets, underscoring its overall superiority. Pose-based approaches inherently exhibits lower bias and reduced susceptibility to background noise, while simultaneously promoting greater privacy and adhering to ethical standards [11, 87, 88].
VI Ablation Study
VI-A Impact of Relative Pose
The results detailed in Table IV empirically validate the effectiveness of using relative pose, as theoretically outlined in Section III-A. This empirical evidence complements and reinforces the theory that incorporating relative movement benefits anomaly detection. The results indicate a consistent improvement across all variations of the SPARTA when the relative pose is integrated. Specifically, regarding AUC-ROC, the SPARTA-H variant incorporating relative movement demonstrates a notable performance enhancement. On the SHT dataset, incorporating relative movement in SPARTA-H improves AUC-ROC by $1.55\%$, demonstrating its effectiveness in capturing motion dynamics. This effect is even more pronounced on CHAD, with a $9.09\%$ improvement, highlighting the significance of relative pose in complex scenarios. On NWPUC [31], SPARTA-C and SPARTA-H exhibit similar gains, while SPARTA-F shows a slight decrease, suggesting dataset-specific variations. Overall, these results affirm that relative movement generally provides complementary information and enhances anomaly detection performance.
TABLE V: Evaluating AUC-ROC on SHT [29], HR-SHT [14], CHAD [30], and NWPUC [31] datasets: A comparative analysis of our design with different methods of tokenization. The best result of each branch is highlighted in gray.
@@ VI-A Impact of Relative Pose
NWPUC
SPARTA-C T-PRP 83.85 85.46 65.59 59.93
KS-PRP 83.67 85.15 64.69 62.67
FS-PRP 83.45 85.13 62.84 61.96
ST-PRP 85.10 86.70 66.12 62.69
SPARTA-F T-PRP 82.08 84.47 66.18 61.85
KS-PRP 79.53 80.40 64.95 61.91
FS-PRP 82.08 83.21 64.29 62.00
ST-PRP 83.19 83.70 66.61 62.29
SPARTA-H T-PRP 84.40 86.06 67.06 61.95
KS-PRP 83.63 84.94 65.37 62.67
FS-PRP 84.68 86.37 64.28 62.45
ST-PRP 85.75 87.23 67.04 63.48
Figure 7: Box-and-whisker plots of tokenization methods’ AUC-ROC performance on SHT [29], HR-SHT [14], CHAD [30], and NWPUC [31] datasets.
VI-B Exploring Diverse Tokenization Strategies
@@
On top of all these tokenization strategies, we also add positional encoding to embed order into input sequences. In Table V, the performance of various tokenization methods is compared. The ST-PRP tokenization method, detailed in Section III-A, demonstrates superior performance in most cases compared to other approaches. While the T-PRP tokenization outperforms others in two specific instances, purely spatial tokenizations (KS-PRP and FS-PRP) consistently yield suboptimal results. Figure 7 further supports this observation, clearly demonstrating that across all four benchmark datasets, the SPARTA variants achieve higher overall AUC-ROC when using ST-PRP tokenization. This suggests that the combination of temporal and spatial attention between tokens uncovers crucial information for analyzing human behavior patterns that neither can detect independently. Consequently, the ST-PRP tokenization method, which integrates spatial and temporal information, emerges as the most effective approach.
VII Conclusion
@@
In this paper, we introduced methodologies that pave the way for advanced human-centric VAD. The novel proposed Spatio-Temporal Pose and Relative Pose (ST-PRP) tokenization method, which serves as a key component for high-level human behavior analysis. Combined with our new Unified Encoder Twin Decoders (UETD) transformer core, the proposed SPARTA architecture demonstrates superior performance in self-supervised human-centric VAD. Extensive benchmarking against SOTA methods confirms SPARTA’s accuracy and robustness. We hope that our contributions will serve as a foundation for future advancements in the field.
Acknowledgments
This research is supported by the National Science Foundation (NSF) under Award Numbers 1831795 and 2329816.
Appendix: Ablation Studies Hyperparameters
@@ Acknowledgments
This section provides an in-depth exposition of the architectural and training hyperparameters of the ablation studies (Section VI) to ensure the reproducibility of the results.
-A Relative Pose Ablation Setup and Hyperparameters
@@
To reveal the benefit of incorporating relative movement, we use the best SPARTA model which includes 12 heads and 4 layers with the feed-forward layer size set to 64. This variant is trained with and without relative movement data using the hyperparameters shown in Table VI. The optimal hyperparameters are chosen using a systematic grid search. In all training instances, we have used the Adam optimizer with a weight decay of $55.0e-$ 5 and trained branches for 30 epochs. For the training, the same strategy is used as in Section III; first, the SPARTA-C (the unified encoder and CTD decoder) is trained. In the next step, the unified encoder is frozen and the FTD decoder is trained. Since both the SHT and HR-SHT utilize identical videos in their training sets, the hyperparameters for both models remain consistent.
@@
TABLE VI: The training hyperparameters used for the Relative Pose Ablation Study (Section VI-A) on SHT [29], HR-SHT [14], CHAD [30], and NWPUC [31] datasets. LR and DR refer to the Learning Rate and Dropout Rate.
@@
LR DR LR DR LR DR
SPARTA-C $32.0e-$ 3 $0.1$ $65.0e-$ 6 $0.1$ $21.0e-$ 2 $0.2$
$51.0e-$ 5 $0.1$ $32.0e-$ 3 $0.1$ $35.0e-$ 3 $0.3$
SPARTA-F $45.0e-$ 4 $0.1$ $33.0e-$ 3 $0.1$ $31.0e-$ 3 $0.1$
$32.0e-$ 3 $0.1$ $45.0e-$ 4 $0.1$ $31.0e-$ 3 $0.1$
Appendix: Tokeniation Ablation Setup and Hyperparamters
To effectively compare various tokenization setups at their full potential, we carried out a systematic grid search. This approach was not just to identify the optimal training hyperparameters, but also to determine the best architectural choices for each setup.
TABLE VII: The design choices used for the Tokenization Ablation Study (Section VI-B)
@@
TABLE VIII: The training hyperparameters used for the Tokenization Ablation Study (Section VI-B) on SHT [29], HR-SHT [14], CHAD [30], and NWPUC [31] datasets. LR and DR refer to the Learning Rate and Dropout Rate.
@@
CHAD NWPUC
LR DR LR DR LR DR
SPARTA-C T-PRP $65.0e-$ 6 $0.1$ $33.0e-$ 3 $0.1$ $31.0e-$ 3 $0.1$
KS-PRP $65.0e-$ 6 $0.1$ $33.0e-$ 3 $0.2$ $31.0e-$ 3 $0.1$
FS-PRP $51.0e-$ 5 $0.1$ $31.0e-$ 3 $0.1$ $31.0e-$ 3 $0.1$
ST-PRP $51.0e-$ 5 $0.1$ $32.0e-$ 3 $0.1$ $35.0e-$ 3 $0.1$
SPARTA-F T-PRP $33.0e-$ 3 $0.1$ $41.0e-$ 4 $0.1$ $32.0e-$ 3 $0.1$
KS-PRP $41.0e-$ 4 $0.2$ $41.0e-$ 4 $0.1$ $33.0e-$ 3 $0.1$
FS-PRP $45.0e-$ 4 $0.1$ $51.0e-$ 5 $0.1$ $31.0e-$ 3 $0.1$
ST-PRP $32.0e-$ 3 $0.1$ $45.0e-$ 4 $0.1$ $31.0e-$ 3 $0.1$
Table VII presents the architectural design parameter choices obtained from the grid search on SHT dataset [29]. These parameters are kept the same for other datasets to ensure a fair comparison. Across all experiments, both the unified encoder and the twin decoders consistently adhere to the parameters detailed in Table VII. This consistency ensures a standardized approach in our experimental setup. On the other hand, Table VIII shows the best learning rate and dropout for training per branch. The number of epochs, optimizer, and weight decay are 30, Adam, and $55.0e-$ 5 respectively.
Regarding the training process, same as other tests, SPARTA-C is initially trained following the strategy outlined in section III, where both the unified encoder and the CTD decoder are trained together. Subsequently, for SPARTA-F, the unified encoder is frozen, and then only the FTD decoder undergoes further training. As the training sets for both the SHT and HR-SHT consist of the same videos, the hyperparameters are kept uniform for them.